OpenStudy (anonymous):

It takes eight hours for a 600 km journey, if 120 km is done by train and the rest by car. It takes 20 minutes more, if 200 km is done by train and the rest by car. The ratio of the speed of the train to that of the cars is:

OpenStudy (anonymous):

the question is worded very badly

OpenStudy (anonymous):

so distance = speed x time

OpenStudy (anonymous):

time = distance / speed

OpenStudy (anonymous):

let the train speed be x, and the car speed be y

OpenStudy (anonymous):

The first sentence tells us that (120/x) +(480/y) = 8 { note: all values are in km and hours , therefore when we solve the equations the units of x and y will be km /hr ( not that it matters because they want the ratio

OpenStudy (anonymous):

the second sentence tells us that (200 / x ) + 400/ y = 8 1/3 ( note 1/3rd of an hour = 20mintues

OpenStudy (anonymous):

then solve them